The Data of Macraconomcs Wark IOut Question 2 Consider an economy that produces and consumes bread and automobiles. The table includes data for two different years 2010 2010 2015 2015 QuantityQuantity Price $65000 Automobiles70 AS $44000110 Bread 530000$10 430000 $19 Round answers to to places after the decimal where necessary b. Calculate the percentage change in each of the following between 2010 and 2015. % change auto prices- % %change bread prices- % change GDP deflator- % %change CPI-

Need help, please show work
Answer #1
Goods 2010
Quantity
2010
Price in $
2015
Quantity
2015
Price in $
Automobiles 70 44000 110 65000
Bread 530000 10 430000 19
(A) Percentage change in Auto Price
Percentage change in Auto Price = (Price of Auto in 2015 - Price of Auto in 2010) / Price of auto in 2010 x 100
Percentage change in Auto Price = (110 - 70) / 70 x 100
Percentage change in Auto Price = 57.14
there is an increase in price of Automobiles by 57.14% from 2010 to 2015
(B) Percentage change in Bread Price
Percentage change in Auto Price = (Price of Bread in 2015 - Price of Bread in 2010) / Price of Bread in 2010 x 100
Percentage change in Auto Price = (430000 - 530000) / 530000 x 100
Percentage change in Auto Price = -18.87
there is decrease in price of bread by 18.87% from 2010 to 2015
(C) Percentage Change in GDP Deflator
Goods 2010
Quantity
2010
Price in $
Expenditure
=Price of 2010 x Quantity of 2010
Automobiles 70 44000 3080000
Bread 530000 10 5300000
8380000
Nominar and Real GDP of 2010 is 8680000, though 2010 is considered as base year
So GDP Deflator of 2010 (Base Year) = Nominal GDP 2010 / Real GDP of 2010 x 100 = 100%
Goods 2015
Quantity
2015
Price in $
Expenditure
=Price of 2015 x Quantity of 2015
2010
Price in $
Expenditure
=Price of 2010 x Quantity of 2015
Automobiles 110 65000 7150000 44000 4840000
Bread 430000 19 8170000 10 4300000
15320000 9140000
Nominal GDP of 2015 = 15320000 and Real GDP of 2015 = 9140000
So GDP Deflator of 2015= Nominal GDP of 2015 / Real GDP of 2015 x 100 = 100%
So GDP Deflator of 2015= 15320000 / 9140000 x 100 = 167.62%
Percentage change in GDP Deflator = (GDP Deflator of 2015 - GDP Deflator of 2010) / GDP Deflator of 2010 x 100
Percentage change in GDP Deflator = (167.62 - 100) / 100 x 100
Percentage change in GDP Deflator = 67.62 %
(D)) Percentage change in CPI
Goods 2010
Quantity
2010
Price in $
Expenditure
=Price of 2010 x Quantity of 2010
2015
Price in $
Expenditure
=Price of 2015 x Quantity of 2010
Automobiles 70 44000 3080000 65000 4550000
Bread 530000 10 5300000 19 10070000
8380000 14620000
CPI value for the year 2010 which is base year in this problem is = 100%
CPI Value for 2015 = Base year quantity multiplied by Price of 2015 / Base year quantity multiplied by Price of 2010 x 100
CPI Value for 2015 = (14620000 / 8380000) x 100
CPI Value for 2015 = 174.46%
Percentage change in CPI = (CPI of 2015 - CPI of 2010) / CPI of 2010 x 100
Percentage change in CPI = (174.46 - 100) / 100 x 100
Percentage change in CPI = 74.46 %